# DiarizationLM: Speaker Diarization Post-Processing with Large Language Models

### v5

@@ 1 Introduction
In many cases, such errors can usually be fixed by leveraging semantic information from the ASR transcripts. Take Fig. 1 as an example, simply by looking at the textual transcript “good morning how are you”, if we know it consists of two speakers, we can easily tell which word comes from which speaker confidently without using any acoustic speaker diarization system. In practice, diarization errors can be much more complicated than the simple example in Fig. 1. To handle such cases, we propose DiarizationLM, a framework to post-process the orchestrated ASR and speaker diarization outputs with a large language model (LLM). While the experiments performed in this paper mainly focus on reducing word diarization errors using LLM, we also show examples how LLM can be used for other purposes such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module in Appendix A.
@@ 2 Motivating example
Apart from correcting diarization errors, we also found several other diarization capabilities of LLMs, such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module. Examples of these capabilities are shown in Appendix A. However, due to lack of relevant training and testing datasets, these capabilities are not thoroughly studied or evaluated in this paper. Here we focus on LLM’s capability of reducing diarization errors.
@@ 3.4 Transcript-Preserving Speaker Transfer
    The transfered speaker label sequence has a 1-to-1 association with the target word sequence.
@@
    The transfered speaker labels are more consistent with the source speaker labels.
As an example, the concatenated word sequence from the completion parser module may not be identical to the original word sequence from the ASR model. Thus we can treat the completion sequences as the source, and the original sequences from the orchestration module as the target, and transfer the speaker labels. Finally, the DiarizationLM outputs will be the original word sequence, associated with the transfered speaker label sequence.
@@
           Transfered speaker sequence of length $M$: $\mathbf{s}^{tra}=(s^{tra}_{1},\cdots,s^{tra}_{M})$
@@ 3.5 LLM finetuning
Although the examples shown in Section 2 and Appendix A were using off-the-shelf Web APIs of commercial LLMs, finetuning the LLM specifically on the speaker diarization task is still required if we need to:
@@ 3.5.1 Flavor 1: hyp2ora
where the output $\mathbf{s}^{ora}$ is the oracle hypothesis speakers transfered from the reference sequences. With $\mathbf{s}^{ora}$, the prompts and completions in our training data are created as below:
@@ 3.5.2 Flavor 2: deg2ref
where the output $\mathbf{s}^{deg}$ is the degraded reference speakers transfered from the hypothesis sequences. With $\mathbf{s}^{deg}$, the prompts and completions in our training data are created as below:
@@ 4.3 Models
For the ASR model in Fig. 2, we use a universal speech model (USM)  [44] with 600 million parameters trained with the RNN-T loss [45]. For the speaker diarization model in Fig. 2, we use the turn-to-diarize system [7] with a multi-stage clustering setup [17] in our experiments, which is capable of diarizing hours of audio recordings in real time on a mobile device [24]. The number of speakers is unknown (from $1$ to $\infty$) to the speaker diarization system in all of our experiments. However, we would like to point out that the proposed framework is very generic and should work with other ASR or speaker diarization systems as well, such as variants of end-to-end speaker diarization models [20, 21, 22, 23].
For the LLM in Fig. 2, we experiment with the PaLM 2-S model (“text-bison” model in Google Cloud API) and the PaLM 2-L model (“text-unicorn” model in Google Cloud API) [46]. We use the PaLM 2-S model as our foundation model, and finetune it on the dataset described in Section 4.1 with data processing steps described in Section 3.5. This model uses a sentence piece model (SPM) of 256k tokens as its tokenizer [47]. During finetuning, we limit the LLM input size by 4,096 tokens, and segment our training and testing data accordingly. The PaLM 2-L model will only be used for zero-shot and one-shot experiments, as described in Section 4.4.
@@ 4.5 Evaluation results
For zero-shot and one-shot experiments with PaLM 2-S, we observe significantly worse WDER and cpWER performance compared with the baseline system, indicating the PaLM 2-S foundation model does not offer speaker diarization capabilities without finetuning. Zero-shot experiment with PaLM 2-L model also shows bad performance, while one-shot experiment with PaLM 2-L model is much better, but still worse than the baseline system. Our results indicate that the PaLM 2-L model with one-shot is able to improve speaker diarization in relatively simple cases as shown in Section 2 and Appendix A. However, real world applications can be much more complicated with errors from both the ASR system and the speaker diarization system. In such cases, even with one-shot, LLM can still introduce even more errors to the results if not finetuned specifically on the speaker diarization task.
@@ 4.6 Case studies
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. WERs are the same for all systems due to TPST. All numbers are percentages.
@@ 5 Discussion and future work
In Appendix A, we have demonstrated other diarization capabilities of LLMs. However, due to lack of relevant datasets, we haven’t been able to thoroughly evaluate these capabilities. One interesting future work would be to collect datasets of these tasks and evaluate how LLM performs.
@@ 7 Conclusion
Appendix A Other diarization capabilities of LLMs
@@
A.1 Autofilling speaker names
@@
A.2 Autofilling speaker roles
@@
A.3 Replacing the orchestration module

### v7

@@ Abstract
Code: https://github.com/google/speaker-id/tree/master/DiarizationLM
Model: https://huggingface.co/google/DiarizationLM-13b-Fisher-v1
@@ 1 Introduction
In many cases, such errors can usually be fixed by leveraging semantic information from the ASR transcripts. Take Fig. 1 as an example, simply by looking at the textual transcript “good morning how are you”, if we know it consists of two speakers, we can easily tell which word comes from which speaker confidently without using any acoustic speaker diarization system. In practice, diarization errors can be much more complicated than the simple example in Fig. 1. To handle such cases, we propose DiarizationLM, a framework to post-process the orchestrated ASR and speaker diarization outputs with a large language model (LLM). While the experiments performed in this paper mainly focus on reducing word diarization errors using LLM, we also show examples how LLM can be used for other purposes such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module in Appendix B.
@@ 2 Motivating example
Apart from correcting diarization errors, we also found several other diarization capabilities of LLMs, such as autofilling speaker names, autofilling speaker roles, or even completely replacing the orchestration module. Examples of these capabilities are shown in Appendix B. However, due to lack of relevant training and testing datasets, these capabilities are not thoroughly studied or evaluated in this paper. Here we focus on LLM’s capability of reducing diarization errors.
@@ 3.4 Transcript-Preserving Speaker Transfer
    The transferred speaker label sequence has a 1-to-1 association with the target word sequence.
@@
    The transferred speaker labels are more consistent with the source speaker labels.
As an example, the concatenated word sequence from the completion parser module may not be identical to the original word sequence from the ASR model. Thus we can treat the completion sequences as the source, and the original sequences from the orchestration module as the target, and transfer the speaker labels. Finally, the DiarizationLM outputs will be the original word sequence, associated with the transferred speaker label sequence.
@@
           transferred speaker sequence of length $M$: $\mathbf{s}^{tra}=(s^{tra}_{1},\cdots,s^{tra}_{M})$
@@ 3.5 LLM finetuning
Although the examples shown in Section 2 and Appendix B were using off-the-shelf Web APIs of commercial LLMs, finetuning the LLM specifically on the speaker diarization task is still required if we need to:
@@ 3.5.1 Flavor 1: hyp2ora
where the output $\mathbf{s}^{ora}$ is the oracle hypothesis speakers transferred from the reference sequences. With $\mathbf{s}^{ora}$, the prompts and completions in our training data are created as below:
@@ 3.5.2 Flavor 2: deg2ref
where the output $\mathbf{s}^{deg}$ is the degraded reference speakers transferred from the hypothesis sequences. With $\mathbf{s}^{deg}$, the prompts and completions in our training data are created as below:
@@ 4.3 Models
For the ASR model in Fig. 2, we use a universal speech model (USM)  [44] with 600 million parameters trained with the RNN-T loss [45]. For the speaker diarization model in Fig. 2, we use the turn-to-diarize system [7] with a multi-stage clustering setup [17] in our experiments, which is capable of diarizing hours of audio recordings in real time on a mobile device [24]. The number of speakers is unknown (from $1$ to $\infty$) to the speaker diarization system in all of our experiments. Specifically, for the Fisher training set, the Fisher testing set, and the Callhome testing set, all utterances have a ground truth number of speakers equal to two, but our turn-to-diarize system may predict any number of hypothesis speakers. The histogram of number of hypothesis speakers predicted by turn-to-diarize on the Fisher training set is shown in Figure 3. Although our experiments are based on the USM + turn-to-diarize setup, we would like to point out that the proposed framework is very generic and should work with other ASR or speaker diarization systems as well, such as variants of end-to-end speaker diarization models [20, 21, 22, 23].
Figure 3: The histogram for the number of hypothesis speakers predicted by the turn-to-diarize system on the Fisher training set. Note that the ground truth number of speakers is always two on the Fisher dataset, but we do not constrain the number of speakers for the turn-to-diarize system.
For the LLM in Fig. 2, we experiment with the PaLM 2-S model (“text-bison” model in Google Cloud API) and the PaLM 2-L model (“text-unicorn” model in Google Cloud API) [46]. We use the PaLM 2-S model as our foundation model, and finetune it on the dataset described in Section 4.1 with data processing steps described in Section 3.5. This model uses a sentence piece model (SPM) of 256k tokens as its tokenizer [47]. During finetuning, we limit the LLM input size by 4,096 tokens, and segment our training and testing data accordingly. The PaLM 2-L model will only be used for zero-shot and one-shot experiments, as described in Section 4.4. We also experimented with open source models such as Llama 2 [48]. Model details on how we finetune these models are provided in Appendix A.
@@ 4.5 Evaluation results
For zero-shot and one-shot experiments with PaLM 2-S, we observe significantly worse WDER and cpWER performance compared with the baseline system, indicating the PaLM 2-S foundation model does not offer speaker diarization capabilities without finetuning. Zero-shot experiment with PaLM 2-L model also shows bad performance, while one-shot experiment with PaLM 2-L model is much better, but still worse than the baseline system. Our results indicate that the PaLM 2-L model with one-shot is able to improve speaker diarization in relatively simple cases as shown in Section 2 and Appendix B. However, real world applications can be much more complicated with errors from both the ASR system and the speaker diarization system. In such cases, even with one-shot, LLM can still introduce even more errors to the results if not finetuned specifically on the speaker diarization task.
@@ 4.6 Case studies
Table 1: Evaluation results of the USM + turn-to-diarize baseline system and the results post-processed by DiarizationLM. For DiarizationLM, we experiment with PaLM 2 foundation models with and without finetuning on the diarization task. We also experiment with Llama 2 models finetuned on the diarization task. WERs are the same for all systems due to TPST. All numbers are percentages.
@@ 5 Discussion and future work
In Appendix B, we have demonstrated other diarization capabilities of LLMs. However, due to lack of relevant datasets, we haven’t been able to thoroughly evaluate these capabilities. One interesting future work would be to collect datasets of these tasks and evaluate how LLM performs.
@@ 7 Conclusion
Appendix A Open source models
Apart from PaLM 2 [46] models, we have also experimented with open source models such as Llama 2 [48]. These finetuned models are publicly available on Hugging Face.
A.1 DiarizationLM-13b-Fisher-v1
This model is finetuned on the training subset of the Fisher corpus, using a LoRA adapter [63] of rank 256. The total number of training parameters is 1,001,390,080. With a batch size of 16, this model has been trained for 12000 steps, which is $\sim$4 epochs of the training data.
We use the “mixed” flavor during our training, meaning we combine data from “hyp2ora” and “deg2ref” flavors (see Section 3.5 for the definition of these flavors). After the prompt builder, we have a total of 48,142 prompt-completion pairs in our training set.
The finetuning took more than 3 days on a Google Cloud VM instance that has one NVIDIA A100 GPU with 80GB memory.
The maximal length of the prompt to this model is 6000 characters, including the “␣-->␣” suffix. The maximal sequence length is 4096 tokens.
The evaluation results of this model is also provided in Table 1. We can see that similar to the finetuned PaLM 2-S models, this model also significantly reduces the WDER and cpWER on the Fisher testing set. However, on the Callhome testing set, we are seeing slight degradation in WDER and cpWER, indicating potential overfitting due to training for $\sim$4 epochs.
Appendix B Other diarization capabilities of LLMs
@@
B.1 Autofilling speaker names
@@
B.2 Autofilling speaker roles
@@
B.3 Replacing the orchestration module